# A Comprehensive Survey on AI-based Methods for Patents

### v2

@@ Abstract
Recent advancements in Artificial Intelligence (AI) and machine learning have demonstrated transformative capabilities across diverse domains. This progress extends to the field of patent analysis and innovation, where AI-based tools present opportunities to streamline and enhance important tasks in the patent cycle such as classification, retrieval, and valuation prediction. This not only accelerates the efficiency of patent researchers and applicants but also opens new avenues for technological innovation and discovery. Our survey provides a comprehensive summary of recent AI tools in patent analysis from more than 40 papers from 26 venues between 2017 and 2023. Unlike existing surveys, we include methods that work for patent image and text data. Furthermore, we introduce a novel taxonomy for the categorization based on the tasks in the patent life cycle as well as the specifics of the AI methods. This interdisciplinary survey aims to serve as a resource for researchers and practitioners who are working at the intersection of AI and patent analysis as well as the patent offices that are aiming to build efficient patent systems.
@@ 1 Introduction
Recent progress in AI and machine learning has shown transformative capabilities across various domains including NLP, computer vision, and healthcare [13, 50]. The field of patents and technological innovation is not an exception. AI-based tools can streamline the complex patent related tasks such as classification, retrieval, and valuation prediction. For instance, for patent examination, patent offices often rely only on the examiner to judge whether a technology is innovative enough and thus, patentable. However, it is challenging for the human examiner to stay updated on various domains due to the exponential growth in technology and apply the knowledge during evaluation. This intersection of AI and patent processes can accelerate the efficiency of the patent systems—patent reviewers as well as applicants—and help in a faster technological innovation to benefit our society.
@@
The patent application and granting process in the patent life cycle involves complex tasks that require significant human effort for both applicants and reviewers. To streamline these complex processes, AI can be helpful, particularly in patent classification, retrieval, and quality analysis [34]. Patent classification can benefit from the AI-based multi-label classification tools for the hierarchical schemes: International Patent Classification (IPC) and the Cooperative Patent Classification [55, 2]. To evaluate novelty and avoid infringement, the patent retrieval task becomes important while filing or reviewing a new patent application. On the other hand, quality analysis also requires a substantial amount of effort. AI-based representation learning methods can be useful in both tasks [10, 43]. Lastly, recent generative AI tools can generate accurate and technical language descriptions for patents and thus, are useful to optimize human resources and precision in patent writing [39].
In the literature, there is a lack of recent surveys on AI tools for patent analysis. The most recent survey [34] does not cover the recent studies in this area. Moreover, it focuses heavily on text-based approaches. Our survey aims to bridge this gap by providing a comprehensive and detailed summary of existing AI methods in more than 40 papers that appeared in 26 different venues from 2017 to 2023 for patent analysis. We include recent AI tools both for images and text data in patent analysis. Moreover, we introduce a novel taxonomy to categorize these methods based on the relevant tasks and the nature of the methods.
1.1 Overview
Figure 1 provides the hierarchical organization where the most important tasks and their related methods are presented. We organize the survey into the following sections: Section 2 provides background on both the relevant tasks in the patent life cycle as well as the patent datasets. Section 3 summarizes the methods for four individual tasks: patent classification, retrieval, quality analysis, and generation. These methods are further grouped based on their commonality. Finally, Section 4 provides a few important research directions including the use of generative AI and multimodal learning. Note, the frequently used AI methods in the papers covered by this survey are referred in Table 2.
@@ 1.1 Overview
Figure 1: The schema of the main organization with a few examples of the methods in each patent related tasks. We summarize the methods for four individual tasks: patent classification, retrieval, quality analysis, and generation. “NN” and “LLM” denote neural network and large language model respectively.
@@ 2 Background
Figure 2: The overview of four major tasks of patent analysis. The patent retrieval task includes obtaining relevant patents (text and images). Please refer to the detailed descriptions of these tasks in Section 2.1.
@@
A patent grants the owner or holder exclusive rights to an invention, and can be a novel product or a process that usually offers a unique method or technical solution. In exchange for this right, inventors must publicly disclose detailed information about their invention in a patent application [^1] https://www.wipo.int/classifications/ipc/. The United States Patent and Trademark Office (USPTO [^2] https://www.uspto.gov/) issues three types of patents: utility, design, and plant. Utility patents protect the rights related to how the invention works or is used. It provides the entitlement to the functionality of a product. On the other hand, design patents protect the right of the look of an invention and are intended to safeguard the form of a product. A plant patent is issued to an inventor who has discovered and invented a unique variety of a plant and asexually reproduced it. In this work, we focus on utility and design patents considering the relevance of the AI tools on the tasks related to these categories. Here, we describe the relevant tasks, the datasets, and their sources.
2.1 Tasks
Patent application and granting processes are complex and involve many complex tasks that include both applicants and patent reviewers. The AI tools can help in simplifying these tasks. Here, we outline the most relevant tasks that can exploit AI-based methods.
2.1.1 Patent Classification
@@
Patent classification is an important and time-consuming task in the patent life cycle [19, 57, 53]. This task involves a multi-label classification for patents where the classification scheme is hierarchical and a patent can get multiple labels. There are two widely used patent classification systems: the International Patent Classification (IPC) and the Cooperative Patent Classification (CPC). The IPC comprises 8 sections, 132 classes, 651 subclasses, 7590 groups, and 70788 subgroups in a hierarchical order (i.e., sections have classes and classes have subclasses, and so on). CPC is an expanded version of the IPC and is collaboratively administered by the European Patent Office (EPO) and the USPTO. It consists of around 250,000 classification entries and is divided into nine sections (A-H and Y), which are further broken down into classes, subclasses, groups, and subgroups [^3] https://www.cooperativepatentclassification.org/. Table 1 shows an example of CPC classification.
Formulation. Given patents as $(x_{i},y_{ik}),i=1,\ldots,N$, where $x_{i}$ are the features of the example patent $i$, $k\in C$, $C$ denotes the set of classes, $y_{ik}\in\{0,1\}$ is a binary indicator of whether class $k$ is the correct classification for the example patent $i$, and $i$ can belong to more than one class in $C$. The goal is to predict $y_{ik}$.

### v3

# A Survey on Patent Analysis: From NLP to Multimodal AI

@@ Abstract
Recent advances in Pretrained Language Models (PLMs) and Large Language Models (LLMs) have demonstrated transformative capabilities across diverse domains. The field of patent analysis and innovation is not an exception, where natural language processing (NLP) techniques presents opportunities to streamline and enhance important tasks—such as patent classification and patent retrieval—in the patent cycle. This not only accelerates the efficiency of patent researchers and applicants, but also opens new avenues for technological innovation and discovery. Our survey provides a comprehensive summary of recent NLP-based methods—including multimodal ones—in patent analysis. We also introduce a novel taxonomy for categorization based on tasks in the patent life cycle, as well as the specifics of the methods. This interdisciplinary survey aims to serve as a comprehensive resource for researchers and practitioners who work at the intersection of NLP, Multimodal AI, and patent analysis, as well as patent offices to build efficient patent systems.
A Survey on Patent Analysis: From NLP to Multimodal AI
Homaira Huda Shomee   Zhu Wang   Sathya N. Ravi   Sourav Medya Department of Computer Science, University of Illinois Chicago {hshome2,zwang260,sathya,medya}@uic.edu
@@ 1 Introduction
The growing complexity and volume of textual data across various domains have driven significant advancements in NLP, particularly through PLMs Devlin et al. (2019) and LLMs Radford et al. (2019). The field of patents and technological innovation is not an exception. This advancement can streamline complex patent-related tasks such as classification, retrieval, and valuation prediction. For instance, for patent examination, patent offices often rely only on the examiner to judge whether a technology is innovative enough and, thus, patentable. However, it is challenging for the human examiner to stay updated on various domains due to the exponential growth in technology and apply the knowledge during evaluation. This intersection of NLP, Multimodal AI, and patent processes can accelerate the efficiency of the patent systems—patent reviewers as well as applicants—and help in a faster technological innovation to benefit our society.
@@
The patent application and granting process involves complex textual analysis tasks that require significant human effort for both applicants and reviewers. To streamline this, NLP techniques can be helpful, particularly in patent classification, retrieval, and quality analysis Krestel et al. (2021). Patent classification can benefit from multi-label classification tools for the hierarchical schemes: International Patent Classification (IPC) and the Cooperative Patent Classification Roudsari et al. (2022); Althammer et al. (2021). To evaluate novelty and avoid infringement, the patent retrieval task becomes important while filing or reviewing a new patent application. On the other hand, quality analysis also requires a substantial amount of effort. NLP-based representation learning methods can be useful in both tasks Chung and Sohn (2020); Lin et al. (2018). Lastly, recent advanced LLMs can generate accurate and technical language descriptions for patents and, thus, are useful to optimize human resources and precision in patent writing Lee and Hsiang (2020a).
The existing patent surveys in the literature Gomez and Moens (2014); Ali et al. (2024); Krestel et al. (2021); Hanbury et al. (2011); Casola and Lavelli (2022) do not cover the recent studies in this area and fail to show the trends and methods in task specific manner. We introduce a novel taxonomy to categorize the methods based on the relevant tasks and the nature of the methods. Our taxonomy provides an in-depth view of the methods being used in specific tasks. Moreover, it captures the recent trends of using advanced methods (e.g., LLMs) that are missing from the existing surveys. This will be beneficial for researchers who aim to build task-specific methods.
Overview. Fig. 1 provides the hierarchical organization of patent tasks and methods. We organize the survey as follows: Sec. 2 provides background, Sec. 3 summarizes the methods for individual tasks, and Sec. 4 provides future research directions. We maintain a GitHub repository for this survey at AI4Patents-survey, which includes categorized papers and other relevant resources.
@@
Figure 1: The schema of the main organization with the methods in each patent-related task. We summarize the methods for four individual tasks: patent classification, retrieval, quality analysis, and generation. “NN”, “MMs”, “PLMs”, and “PatLMs” denote neural networks, multimodal models, pre-trained language models, and patent language models, respectively. The works that use patent images are written in blue.
@@ 2 Background
A patent grants the owner or holder exclusive rights to an invention and can be a novel product or a process that usually offers a unique method or technical solution. In exchange for this right, inventors must publicly disclose detailed information about their invention in a patent application. The United States Patent and Trademark Office (USPTO [^1] https://www.uspto.gov/) issues three types of patents: utility, design, and plant. In this work, we focus on utility and design patents, considering their importance in innovation across industries. Utility patents protect the rights related to how the invention works or is used. It provides the entitlement to the functionality of a product. On the other hand, design patents protect the right of the look of an invention and are intended to safeguard the form of a product. Here, we outline the relevant tasks.
Formulation. We provide the problem formulations of these patent tasks in Appendix A.3.
Datasets. We describe the common benchmark patent datasets in Appendix A.6.
2.1 Patent Classification
@@
Patent classification is an important but time-intensive task in the patent life cycle Grawe et al. (2017); Shalaby et al. (2018); Risch and Krestel (2018). This involves a multi-label classification for patents where the classification scheme is hierarchical, and a patent can get multiple labels. There are two widely used patent classification systems: International Patent Classification (IPC) and the Cooperative Patent Classification (CPC). The IPC comprises 8 sections, 132 classes, 651 subclasses, 7590 groups, and 70788 subgroups in a hierarchical order (i.e., sections have classes and classes have subclasses, and so on). CPC is an expansion of IPC and is collaboratively administered by the European Patent Office (EPO) and the USPTO. It consists of around 250,000 classification entries and is divided into nine sections (A-H and Y), which are further broken down into classes, subclasses, groups, and subgroups [^2] https://www.cooperativepatentclassification.org/. Table 7 (see Appendix) shows an example of CPC classification.